The owner of a coffee shop has found that the average amount spent by a sample of 25 customers is $8.60 with a sample standard deviation of $1.4. What are the lower and upper bounds of a 90% confidence interval for the population mean amount spent (in dollars) by customers? (6.32, 7.28) (6.34, 7.26) (8.12, 9.08) (9.50, 9.70)
Answer #1

Sample size=25

Standard deviation=S.D.= $1.4

Mean= $8.6

Z value =1.64

Marginal of error= Z * S.D/√N= 1.64*(1.4/√25)= 0.47

Lower bound= 8.6-0.47= 8.13

Upper bound=8.6+0.47= 9.07

Answer= (8.12,9.08)